A car of mass m follows a truck of mass 2m around a circular turn. Both vehicles move at speed v.


(a) What is the ratio of the truck's net centripetal force to the car's net centripetal force?

Ftruck

____

Fcar = 


(b) At what new speed vtruck will the net centripetal force acting on the truck equal the net centripetal force acting on the car still moving at the original speed v? State your answer in terms of v.

vtruck =  


1
Expert's answer
2021-10-03T13:22:01-0400

(a) Formula for centripetal force is

"F=\frac{mv^2}{r} \\

F_{truck} = \frac{2mv^2}{r} \\

F_{car} = \frac{mv^2}{r} \\

\frac{F_{truck}}{F_{car}} = 2"

(b)

"F_{car} = \frac{mv^2}{r} \\

F_{truck} = \frac{2mv_{truck}^2}{r} \\

F_{car} = F_{truck} \\

\frac{2mv_{truck}^2}{r} = \frac{mv^2}{r} \\

v_{truck} = \sqrt{\frac{v^2}{2}} \\

= \frac{v}{\sqrt{2}}"